# Arctic-Embed 2.0: Multilingual Retrieval Without Compromise

### v1

@@ 1 Introduction
We develop Arctic-Embed 2.0 to deliver frontier retrieval quality while addressing two key limitations observed in current multilingual embedding models:
@@ 2.4 Hard Negative Mining for Finetuning
To select the most effective hard negatives, we adopt the strategy from NV Retriever (Moreira et al., 2024): documents scored as most relevant by a teacher embedding model are used as negatives, but any negative with a relevance score exceeding a specified percentage of the known-positive’s score is discarded as a potential false negative. We used stella-en-1.5B-v5111https://huggingface.co/dunzhang/stella_en_1.5B_v5 as the English teacher model, and multilingual-e5-large for multilingual data. Using gte-large-en-v1.5 for comparison, we confirm Moreira et al. (2024)’s finding that stronger teacher models yield higher-quality fine-tuning datasets (Figure 2, left). Rather than adhering to the $95\%$ false-positive filtering threshold suggested in prior work, however, we experimented with varying thresholds and observed improvement at higher thresholds (Figure 2, middle).
@@ Appendix B Pretraining Data Breakdown
We apply top-20-in-3M-shard retrieval-based consistency filtering to each language subset of each dataset, resulting in approximately 1.13 billion unsupervised query-document pairs. A detailed breakdown of this combined dataset by source and language is shown in Figure 5.
Figure 5: Breakdown of 1.13B contrastive pretraining samples by data source (top) and language (bottom).

### v2

@@ 1 Introduction
We develop Arctic-Embed 2.0111The open-source model weights are available under the Apache 2.0 License: snowflake-arctic-embed-m-v2.0 and snowflake-arctic-embed-l-v2.0. to deliver frontier retrieval quality while addressing two key limitations observed in current multilingual embedding models:
@@ 2.4 Hard Negative Mining for Finetuning
To select the most effective hard negatives, we adopt the strategy from NV Retriever (Moreira et al., 2024): documents scored as most relevant by a teacher embedding model are used as negatives, but any negative with a relevance score exceeding a specified percentage of the known-positive’s score is discarded as a potential false negative. We used stella-en-1.5B-v5222https://huggingface.co/dunzhang/stella_en_1.5B_v5 as the English teacher model, and multilingual-e5-large for multilingual data. Using gte-large-en-v1.5 for comparison, we confirm Moreira et al. (2024)’s finding that stronger teacher models yield higher-quality fine-tuning datasets (Figure 2, left). Rather than adhering to the $95\%$ false-positive filtering threshold suggested in prior work, however, we experimented with varying thresholds and observed improvement at higher thresholds (Figure 2, middle).
@@ Appendix B Pretraining Data Breakdown
We apply top-20-in-3M-shard retrieval-based consistency filtering to each language subset of each dataset, resulting in approximately 1.41 billion unsupervised query-document pairs. A detailed breakdown of this combined dataset by source and language is shown in Figure 5.
Figure 5: Breakdown of 1.41B contrastive pretraining samples by data source (top) and language (bottom).